### v1

# A short proof of generalized Conway–Gordon–Sachs theorem

Abstract.
The famous Conway–Gordon–Sachs theorem for the complete graph on six vertices was extended to the general complete graph on $n$ vertices by Kazakov–Korablev as a congruence modulo $2$, and its integral lift was given by Morishita–Nikkuni. However, the proof is complicated and long. In this paper, we provide a shorter proof of the generalized Conway–Gordon–Sachs theorem over integers.
Key words and phrases:
Spatial graphs, Conway–Gordon–Sachs theorem
1991 Mathematics Subject Classification:
Primary 57M15; Secondary 57K10
The author was supported by JSPS KAKENHI Grant Number JP19K03500.
1. Introduction
An embedding $f$ of a finite graph $G$ into the $3$-sphere is called a spatial embedding of $G$, and $f(G)$ is called a spatial graph of $G$. We call a subgraph of $G$ homeomorphic to the circle a cycle of $G$, and a cycle containing exactly $p$ vertices a $p$-cycle. We denote the set of all $p$-cycles of $G$ by $\Gamma_{p}(G)$. Additionally, we also denote the set of all disjoint pairs of cycles of $G$ consisting of a $p$-cycle and a $q$-cycle by $\Gamma_{p,q}(G)$. In general, for a closed $1$-manifold $\lambda$ in $G$, the image $f(\lambda)$ is no other than a knot or a link contained in $f(G)$. In particular, if $\lambda$ belongs to $\Gamma_{p,q}(G)$, then we call $f(\lambda)$ a $(p,q)$-link of $f(G)$. Moreover, if $\lambda$ contains all vertices of $G$, then we also call $f(\lambda)$ a Hamiltonian link of $f(G)$.
Let $K_{n}$ be the complete graph on $n$ vertices, that is the graph consisting of $n$ vertices such that each pair of two distinct vertices is connected by exactly one edge. The fact that for every spatial graph of $K_{6}$, the sum of the linking numbers over all of the $(3,3)$-links is odd is well known as the Conway-Gordon–Sachs Theorem [1], [7]. If $n$ is 7 or more, Kazakov–Korablev showed that for every spatial graph of $K_{n}$, the sum of the linking numbers over all of the Hamiltonian links is even [2]. On the other hand, Morishita and the author significantly generalized this modulo $2$ congruence formula over integers as follows.
Theorem 1-1.
(Morishita–Nikkuni [4]) Let $n\geq 6$ be an integer and $p,q\geq 3$ two integers satisfying $n=p+q$. For every spatial graph $f(K_{n})$ of $K_{n}$, the following holds:
(1.1) $\sum_{\lambda\in\Gamma_{p,q}(K_{n})}{\rm lk}(f(\lambda))^{2}=(2-\delta_{pq})\cdot(n-6)!\sum_{\lambda\in\Gamma_{3,3}(K_{n})}{\rm lk}(f(\lambda))^{2}.$
Here, lk denotes the linking number and $\delta_{pq}$ denotes the Kronecker’s delta.
It also holds from (1.1) that
(1.2) $\sum_{p+q=n}\sum_{\lambda\in\Gamma_{p,q}(K_{n})}{\rm lk}(f(\lambda))^{2}=(n-5)!\sum_{\lambda\in\Gamma_{3,3}(K_{n})}{\rm lk}(f(\lambda))^{2}.$
Kazakov–Korablev’s result above can be obtained immediately by taking modulo $2$ reduction on both sides of (1.2). Furthermore, since the right side of (1.2) is not zero, it is also clear that for any $p$ and $q$ satisfying $p+q=n$, there always exists a nonsplitable Hamiltonian $(p,q)$-link of $f(K_{n})$. This is also a generalization of Vesnin–Litvintseva’s result [8] that $f(K_{n})$ always contain a nonsplittable Hamiltonian link. For additional formulas on evaluating the sum of linking numbers over Hamiltonian links derived from (1.1), see [4]. The proof of Theorem 1-1 given in [4] was done using tricky induction procedures and was lengthy and tedious. Our purpose in this paper is to simplify and shorten the proof of Theorem 1-1, thereby making it more accessible to a wider reader. The key to this simplification is Lemma 2-2, which reveals a particular homological property concerning the linking of a knot with a spatial complete graph. The proof of Theorem 1-1 presented here is a reorganized version of the one originally introduced by the author in the Japanese textbook [6] several years ago.
Remark 1-2.
Conway–Gordon also showed that for every spatial graph of $K_{7}$, the sum of the Arf invariants over all of the Hamiltonian knots is odd [1]. For further details on refining this fact or extending it to general $K_{n}$, see [5], [3].
2. New proof of Theorem 1-1
In the following, let $K_{n}$ consist of $n$ vertices $1,2,\ldots,n$. We denote the edge of $K_{n}$ connecting two distinct vertices $i$ and $j$ by $\overline{ij}$, and denote a path of length 2 of $K_{n}$ consisting of two edges $\overline{ij}$ and $\overline{jk}$ by $\overline{ijk}$.
Lemma 2-1.
Let $G$ be the graph which is a disjoint union of a loop $e$ and $K_{4}$. For every spatial graph $f(G)$ of $G$, the following holds:
(2.1) $\sum_{\delta\in\Gamma_{4}(K_{4})}{\rm lk}(f(e),f(\delta))^{2}=\sum_{\gamma\in\Gamma_{3}(K_{4})}{\rm lk}(f(e),f(\gamma))^{2}.$
Proof.
We give the orientation to the loop $e$ and each of the edges of $K_{4}$ as shown in Fig. 2.1, and identify each of the cycles of $K_{4}$ with an element of the first integral homology group $H_{1}(K_{4};{\mathbb{Z}})$ as follows:
$\gamma_{1}=\langle 12\rangle+\langle 24\rangle-\langle 14\rangle,\ \gamma_{2}=\langle 23\rangle+\langle 34\rangle-\langle 24\rangle,\ \gamma_{3}=\langle 31\rangle+\langle 14\rangle-\langle 34\rangle,$
$\gamma_{4}=\langle 12\rangle+\langle 23\rangle+\langle 31\rangle,\ \delta_{1}=\langle 23\rangle+\langle 31\rangle+\langle 14\rangle-\langle 24\rangle,$
$\delta_{2}=\langle 31\rangle+\langle 12\rangle+\langle 24\rangle-\langle 34\rangle,\ \delta_{3}=\langle 12\rangle+\langle 23\rangle+\langle 34\rangle-\langle 14\rangle.$
Note that $\gamma_{i}\in\Gamma_{3}(K_{4})\ (i=1,2,3,4)$ and $\delta_{j}\in\Gamma_{4}(K_{4})\ (j=1,2,3)$. Since $\delta_{j}=\gamma_{j+1}+\gamma_{j+2}$ in $H_{1}(K_{4};{\mathbb{Z}})$, we have
${\rm lk}(f(e),f(\delta_{j}))$ $=$ ${\rm lk}(f(e),f(\gamma_{j+1}))+{\rm lk}(f(e),f(\gamma_{j+2})),$
where we regard $\gamma_{j+3}=\gamma_{j}$. Then we have
$\sum_{j=1}^{3}{\rm lk}(f(e),f(\delta_{j}))^{2}$ $=$ $2\sum_{i=1}^{3}{\rm lk}(f(e),f(\gamma_{i}))^{2}$
$+2\sum_{1\leq i<j\leq 3}{\rm lk}(f(e),f(\gamma_{i})){\rm lk}(f(e),f(\gamma_{j})).$
On the other hand, since $\gamma_{4}=\gamma_{1}+\gamma_{2}+\gamma_{3}$ in $H_{1}(K_{4};{\mathbb{Z}})$, we have
(2.3) ${\rm lk}(f(e),f(\gamma_{4}))=\sum_{i=1}^{3}{\rm lk}(f(e),f(\gamma_{i})).$
By (2) and (2.3), we have
$\sum_{i=1}^{4}{\rm lk}(f(e),f(\gamma_{i}))^{2}$ $=$ $\sum_{i=1}^{3}{\rm lk}(f(e),f(\gamma_{i}))^{2}+{\rm lk}(f(e),f(\gamma_{4}))^{2}$
$=$ $2\sum_{i=1}^{3}{\rm lk}(f(e),f(\gamma_{i}))^{2}$
$+2\sum_{1\leq i<j\leq 3}{\rm lk}(f(e),f(\gamma_{i})){\rm lk}(f(e),f(\gamma_{j}))$
$=$ $\sum_{j=1}^{3}{\rm lk}(f(e),f(\delta_{j}))^{2}.$
This completes the proof. ∎
Figure 2.1. Oriented loop $e$ and oriented $K_{4}$
Lemma 2-2.
Let $n\geq 4$ be an integer and $G$ the graph which is a disjoint union of a loop $e$ and $K_{n}$. For every spatial graph $f(G)$ of $G$, the following holds:
(2.4) $\sum_{\gamma\in\Gamma_{n}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}=\sum_{\gamma\in\Gamma_{n-1}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}.$
Proof.
Let us prove (2.4) by induction on $n$. The case where $n=4$ is nothing other than Lemma 2-1. So we assume that $n>4$. Let $f(G)$ be a spatial graph of $G$. For a vertex $m$ of $K_{n}$ $(m=1,2,\ldots,n)$, we denote the subgraph obtained by removing $m$ and all edges incident to $m$ from $K_{n}$ by $K_{n-1}^{(m)}$. This subgraph is isomorphic to $K_{n-1}$, see the left figure in Fig. 2.2. Furthermore, for vertices $i,j$ and $m$ of $K_{n}$ such that $1\leq i<j\leq n$ and $i,j\neq m$, let $F_{ij}^{(m)}$ denote the subgraph obtained by removing the edge $\overline{ij}$ and $n-3$ edges $\overline{mk}$ $(1\leq k\leq n,\ k\neq i,j)$. This subgraph is not isomorphic to $K_{n-1}$ but is homeomorphic to it. Actually, $F_{ij}^{(m)}$ is obtained by replacing the edge $\overline{ij}$ in $K_{n-1}^{(m)}$ with the path $\overline{imj}$ of length $2$, see the right figure in Fig. 2.2. Then for the spatial subgraph $f(e\cup F_{ij}^{(m)})$, by the assumption we have
$\sum_{\gamma\in\Gamma_{n}(F_{ij}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}+\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(K_{n-1}^{(m)})\\ \overline{ij}\not\subset\gamma\end{subarray}}{\rm lk}(f(e),f(\gamma))^{2}$
$=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(F_{ij}^{(m)})\\ \overline{imj}\subset\gamma\end{subarray}}{\rm lk}(f(e),f(\gamma))^{2}+\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-2}(K_{n-1}^{(m)})\\ \overline{ij}\not\subset\gamma\end{subarray}}{\rm lk}(f(e),f(\gamma))^{2}.$
Let us fix the vertex $m$ and add both sides of (2) for all $i,j$ such that $1\leq i<j\leq n$ and $i,j\neq m$. First, consider an $n$-cycle $\gamma$ of $K_{n}$ where the vertices $i$ and $j$ are adjacent to the vertex $m$ on $\gamma$. In this case, $\gamma$ is an $n$-cycle of $F_{ij}^{(m)}$ and we have
(2.6) $\sum_{\begin{subarray}{c}1\leq i<j\leq n\\ i,j\neq m\end{subarray}}\bigg{(}\sum_{\gamma\in\Gamma_{n}(F_{ij}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}=\sum_{\gamma\in\Gamma_{n}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}.$
Next, consider an $(n-1)$-cycle $\gamma$ of $K_{n-1}^{(m)}$. Let $\overline{ij}$ be an edge of $K_{n-1}^{(m)}$ that is not contained in $\gamma$. Then there are $\binom{n-1}{2}-(n-1)=(n^{2}-5n+4)/2$ ways to choose such $i$ and $j$, and thus we have
$\sum_{\begin{subarray}{c}1\leq i<j\leq n\\ i,j\neq m\end{subarray}}\bigg{(}\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(K_{n-1}^{(m)})\\ \overline{ij}\not\subset\gamma\end{subarray}}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}$
$=$ $\frac{n^{2}-5n+4}{2}\sum_{\gamma\in\Gamma_{n-1}(K_{n-1}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}.$
Next, consider an $(n-1)$-cycle $\gamma$ of $K_{n}$ containing the vertex $m$, where the vertices $i$ and $j$ are adjacent to $m$ on $\gamma$. Then $\gamma$ is an $(n-1)$-cycle of $F_{ij}^{(m)}$ containing the path $\overline{imj}$ and we have
(2.8) $\sum_{\begin{subarray}{c}1\leq i<j\leq n\\ i,j\neq m\end{subarray}}\bigg{(}\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(F_{ij}^{(m)})\\ \overline{imj}\subset\gamma\end{subarray}}\!\!\!\!{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}$ $=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(K_{n})\\ m\subset\gamma\end{subarray}}\!\!\!\!{\rm lk}(f(e),f(\gamma))^{2}.$
Finally, consider an $(n-2)$-cycle $\gamma$ of $K_{n-1}^{(m)}$. Let $\overline{ij}$ be an edge of $K_{n-1}^{(m)}$ that is not contained in $\gamma$. Then there are $\binom{n-1}{2}-(n-2)=(n^{2}-5n+6)/2$ ways to choose such $i$ and $j$, and thus we have
$\sum_{\begin{subarray}{c}1\leq i<j\leq n\\ i,j\neq m\end{subarray}}\bigg{(}\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-2}(K_{n-1}^{(m)})\\ \overline{ij}\not\subset\gamma\end{subarray}}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}$
$=$ $\frac{n^{2}-5n+6}{2}\sum_{\gamma\in\Gamma_{n-2}(K_{n-1}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}.$
Furthermore, for each $f(K_{n-1}^{(m)})$, the following holds by the assumption:
(2.10) $\sum_{\gamma\in\Gamma_{n-1}(K_{n-1}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}=\sum_{\gamma\in\Gamma_{n-2}(K_{n-1}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}.$
Theorefore, from (2) and (2.6), (2), (2.8), (2), (2.10), we have
$\sum_{\gamma\in\Gamma_{n}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}$
$=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(K_{n})\\ m\subset\gamma\end{subarray}}{\rm lk}(f(e),f(\gamma))^{2}+\sum_{\gamma\in\Gamma_{n-1}(K_{n-1}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}.$
Then, let us add both sides of (2) for all $m=1,2,\ldots,n$. First, in the $(n-1)$-cycle $\gamma$ of $K_{n}$, there are $(n-1)$ ways to choose the vertex $m$ that is contained in $\gamma$. Thus we have
(2.12) $\sum_{m=1}^{n}\bigg{(}\sum_{\begin{subarray}{c}\gamma\in\Gamma_{n-1}(K_{n})\\ m\subset\gamma\end{subarray}}\!\!\!\!{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}=(n-1)\sum_{\gamma\in\Gamma_{n-1}(K_{n})}\!\!\!\!{\rm lk}(f(e),f(\gamma))^{2}.$
Second, in the $(n-1)$-cycle $\gamma$ of $K_{n}$, for the vertex $m$ that is not contained in $\gamma$, the cycle $\gamma$ is an $(n-1)$-cycle of $K_{n-1}^{(m)}$. Since there is only one way to choose such an $m$, we have
(2.13) $\sum_{m=1}^{n}\bigg{(}\sum_{\gamma\in\Gamma_{n-1}(K_{n-1}^{(m)})}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}=\sum_{\gamma\in\Gamma_{n-1}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}.$
Then by substituting (2.12) and (2.13) into (2), we have
$n\sum_{\gamma\in\Gamma_{n}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}=n\sum_{\gamma\in\Gamma_{n-1}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}.$
This completes the proof. ∎
Figure 2.2. $K_{n-1}^{(m)}$ (left) and $F_{ij}^{(m)}$ (right) $(n=7)$
Corollary 2-3.
Let $n\geq 4$ be an integer and $G$ the graph which is a disjoint union of a loop $e$ and $K_{n}$. For every spatial graph $f(G)$ of $G$, the following holds:
(2.14) $\sum_{\gamma\in\Gamma_{n}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}=(n-3)!\sum_{\gamma\in\Gamma_{3}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}.$
Proof.
We denote the subgraph of $K_{n}$ obtained by removing exactly $k$ vertices $m_{1},m_{2},\ldots,m_{k}\ (1\leq k\leq n-1)$ and all edges incident to them from $K_{n}$ by $K_{n-k}^{(m_{1}m_{2}\cdots m_{k})}$. Note that this subgraph is isomorphic to $K_{n-k}$. Then by repeatedly using Lemma 2-2, we have
$\sum_{\gamma\in\Gamma_{n}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}=\sum_{\gamma\in\Gamma_{n-1}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}$
$=$ $\sum_{m_{1}=1}^{n}\bigg{(}\sum_{\gamma\in\Gamma_{n-1}(K_{n-1}^{(m_{1})})}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}$
$=$ $\sum_{m_{1}=1}^{n}\bigg{(}\sum_{\gamma\in\Gamma_{n-2}(K_{n-1}^{(m_{1})})}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}=2\sum_{\gamma\in\Gamma_{n-2}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}$
$=$ $2\sum_{1\leq m_{1}<m_{2}\leq n}\bigg{(}\sum_{\gamma\in\Gamma_{n-2}(K_{n-2}^{(m_{1}m_{2})})}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}$
$=$ $2\sum_{1\leq m_{1}<m_{2}\leq n}\bigg{(}\sum_{\gamma\in\Gamma_{n-3}(K_{n-2}^{(m_{1}m_{2})})}{\rm lk}(f(e),f(\gamma))^{2}\bigg{)}$
$=$ $2\cdot 3\sum_{\gamma\in\Gamma_{n-3}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}=\cdots=k!\sum_{\gamma\in\Gamma_{n-k}(K_{n})}{\rm lk}(f(e),f(\gamma))^{2}.$
So if we set $k=n-3$, we have the desired conclusion. ∎
Proof of Theorem 1-1.
For a $k$-cycle $\gamma$ of $K_{n}$, we denote the subgraph of $K_{n}$ obtained by removing all $k$ vertices of $\gamma$ and all edges incident to them from $K_{n}$ by $G_{\gamma}$. Note that this subgraph is isomorphic to $K_{n-k}$. Then we have
$(1+\delta_{pq})\sum_{\lambda\in\Gamma_{p,q}(K_{n})}{\rm lk}(f(\lambda))^{2}$
$=$ $\sum_{\gamma\in\Gamma_{p}(K_{n})}\bigg{(}\sum_{\gamma^{\prime}\in\Gamma_{q}(G_{\gamma})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}.$
Then by Corollary 2-3, the right side of (2) can be expressed as follows:
$\sum_{\gamma\in\Gamma_{p}(K_{n})}\bigg{(}\sum_{\gamma^{\prime}\in\Gamma_{q}(G_{\gamma})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}$
$=$ $\sum_{\gamma\in\Gamma_{p}(K_{n})}\bigg{(}(q-3)!\sum_{\gamma^{\prime}\in\Gamma_{3}(G_{\gamma})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}$
$=$ $(q-3)!\sum_{\gamma^{\prime}\in\Gamma_{3}(K_{n})}\bigg{(}\sum_{\gamma\in\Gamma_{p}(G_{\gamma^{\prime}})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}.$
Let $H_{\gamma^{\prime}}^{i}\ (i=1,2,\ldots,l,\ l=\binom{n-3}{p})$ be the collection of all subgraphs of $G_{\gamma^{\prime}}$ that are isomorphic to $K_{p}$. Then the sets of $p$-cycles $\Gamma_{p}(H_{\gamma^{\prime}}^{i})$ are mutually disjoint, and their direct sum equals $\Gamma_{p}(G_{\gamma^{\prime}})$. Then by Corollary 2-3, we have
$\sum_{\gamma^{\prime}\in\Gamma_{3}(K_{n})}\bigg{(}\sum_{\gamma\in\Gamma_{p}(G_{\gamma^{\prime}})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}$
$=$ $\sum_{\gamma^{\prime}\in\Gamma_{3}(K_{n})}\bigg{(}\sum_{i=1}^{l}\bigg{(}\sum_{\gamma\in\Gamma_{p}(H_{\gamma^{\prime}}^{i})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}\bigg{)}$
$=$ $\sum_{\gamma^{\prime}\in\Gamma_{3}(K_{n})}\bigg{(}\sum_{i=1}^{l}\bigg{(}(p-3)!\sum_{\gamma\in\Gamma_{3}(H_{\gamma^{\prime}}^{i})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}\bigg{)}$
$=$ $(p-3)!\sum_{\gamma^{\prime}\in\Gamma_{3}(K_{n})}\bigg{(}\sum_{i=1}^{l}\bigg{(}\sum_{\gamma\in\Gamma_{3}(H_{\gamma^{\prime}}^{i})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}\bigg{)}.$
Note that each $3$-cycle $\gamma$ of $G_{\gamma^{\prime}}$ is shared by exactly $\binom{n-6}{p-3}$ subgraphs $H_{\gamma^{\prime}}^{i}$’s. This implies that
(2.18) $\sum_{i=1}^{l}\bigg{(}\sum_{\gamma\in\Gamma_{3}(H_{\gamma^{\prime}}^{i})}\!\!\!\!{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}=\binom{n-6}{p-3}\sum_{\gamma\in\Gamma_{3}(G_{\gamma^{\prime}})}\!\!\!\!{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}.$
Then it holds from (2), (2) and (2.18) that
$\sum_{\gamma\in\Gamma_{p}(K_{n})}\bigg{(}\sum_{\gamma^{\prime}\in\Gamma_{q}(G_{\gamma})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}$
$=$ $(p-3)!(q-3)!\binom{n-6}{p-3}\sum_{\gamma^{\prime}\in\Gamma_{3}(K_{n})}\bigg{(}\sum_{\gamma\in\Gamma_{3}(G_{\gamma^{\prime}})}{\rm lk}(f(\gamma),f(\gamma^{\prime}))^{2}\bigg{)}$
$=$ $2(n-6)!\sum_{\lambda\in\Gamma_{3,3}(K_{n})}{\rm lk}(f(\lambda))^{2}.$
By (2), (2) and $2/(1+\delta_{pq})=2-\delta_{pq}$, we have the result. ∎